Quiz
ECON 2300 - Chapter 4 Problem Set
Erstelle ein kostenloses Konto, um dieses Quiz zu beenden
Speichere deinen Fortschritt und sieh dir dein Endergebnis an.
oder mit deiner E-Mail registrieren
Schon ein Konto? Anmelden
Fragen in diesem Quiz (23)
1The _____ of an event is a number that measures the likelihood that an event will occur when an experiment is carried out.
- AOutcome
- BProbability
- CIntersection
- DObservation
2Two mutually exclusive events having positive probabilities are ______________ dependent.
- AAlways
- BSometimes
- CNever
3What is the key difference between combinations and permutations?
- ACombinations do not factor in the order of n objects and permutations do
- BCombinations require that n objects are mutually exclusive and permutations do not
- CThe Counting Rule for Combinations formula utilizes factorials and the Counting Rule for Permutations formula does not
4Which of the following is not a basic requirement for assigning probabilities?
- AThe probability assigned to each experimental outcome must be between 0 and 1
- BThe sum of the probabilities for all the experimental outcomes must equal 1.0
- CThe mean of the probabilities for all the experimental outcomes must equal 0.5
5When assigning probability to whether a specific number will appear on the upward face of a standard six-sided die, P(2) would be equal to ________.
- A1/6
- B2/6 or 1/3
- C6/2
- DA very sad amount of damage to a Gelatinous Cube while playing D&D
6___________ of assigning probabilities is most appropriate when one cannot realistically assume that the experimental outcomes are equally likely and when little relevant data is available.
- AThe subjective method
- BThe relative frequency method
- CThe classical method
- DA [meth]od addict
7How many ways can three items be selected from a group of six items?
- A20
- B2
- C0.5
- D40
8What would be the best way to represent the intersection of events A and B?
- AVenn Diagram
- Baddition law
- Csample space
- Dthe classical method
9If two events are independent, we can _____ their probabilities to determine the intersection probability.
- ADivide
- BAdd
- CMultiply
- DSubtract
10What does P(A | B) translate to?
- Athe probability of A given B
- BThe union of A and B
- Cthe sample space outside of A and B
- Dthe probabilities between event A and event B
11A ____________ is the probability that one event will occur given that we know that another event already has occurred.
- Asample space outcome
- Bsubjective probability
- Ccomplement of events
- Dlong-run relative frequency
- Econditional probability
12The _______ of two events X and Y is another event that consists of the sample space outcomes belonging to either event X or event Y or both event X and Y
- Acomplement
- Bunion
- Cintersection
- Dconditional probability
13P(A U B) = P(A) + P(B) - P(A ∩ B) represents the formula for the _____________.
- Aconditional probability
- Baddition law
- Caddition law for two mutually exclusive events
- Dmultiplication rule
14Suppose a certain ophthalmic trait is associated with eye color. Three hundred randomly selected individuals are studied with results as follows:
EYE COLOR
TRAIT | Blue | Brown | Other | Total
____________________________________
Yes | 70 | 30 | 20 | 120
____________________________________
No | 20 | 110 | 50 | 180
____________________________________
Total | 90 | 140 | 70 | 300
Which of the following expressions describes the relationship between the events A (a person has brown eyes) and B (a person has blue eyes)?- Aindependent
- Bexhaustive
- Csimple
- Dmutually exclusive
15Among twenty-five articles, nine are defective, six having only minor defects and three having major defects. Determine the probability that an article selected at random has major defects given that it has defects.
- A1/3
- B1/4
- C.24
- D.08
16The depositors at Save-More Bank are categorized by age and sex. We are going to select an individual at random from this group of 2000 depositors.
Sex
Age | Male | Female
-------------------------------------
30 or less | 800 | 600
31 or more | 400 | 200
-------------------------------------
P[Female(30 or less)] =- A2/5
- B3/4
- C3/7
- D3/10
17The depositors at Save-More Bank are categorized by age and sex. We are going to select an individual at random from this group of 2000 depositors.
Sex
Age | Male | Female
-------------------------------------
30 or less | 800 | 600
31 or more | 400 | 200
-------------------------------------
P[Male or (31 or more)]- A1/5
- B3/10
- C1/2
- D7/10
18The depositors at Save-More Bank are categorized by age and sex. We are going to select an individual at random from this group of 2000 depositors.
Sex
Age | Male | Female
-------------------------------------
30 or less | 800 | 600
31 or more | 400 | 200
-------------------------------------
P(Female) =- A2/5
- B2/3
- C3/4
- D3/5
19The depositors at Save-More Bank are categorized by age and sex. We are going to select an individual at random from this group of 2000 depositors.
Sex
Age | Male | Female
-------------------------------------
30 or less | 800 | 600
31 or more | 400 | 200
-------------------------------------
What is the conditional probability that the depositor drawn is 30 or less, given that he is a male?- A2/3
- B7/10
- C4/7
- D2/5
- ENone of the above
20A dormitory on campus houses 200 students. 120 are male, 50 are upper division students, and 40 are upper division male students. A student is selected at random. The probability of selecting a lower division student, given the student is a female, is:
- A7/8
- B7/15
- C2/5
- D7/20
- E1/4
21A local trade union consists of plumbers and electricians. Classified according to rank:
Apprentice Journeyman Master
------------------------------------
Plumbers 25 | 20 | 30 75
------------------------------------
Electricians 15 | 40 | 20 75
-------------------------------------
40 60 50
A member of the union is selected at random. Given that the person selected is a plumber, the probability that he is a journeyman is:- A1/2
- B1/3
- C4/15
- D2/15
- ENone of the above
22Which of the following is false?
- AAlways P(A∪B)=P(B∪A)
- BAlways P(A∩B)=P(B∩A)
- CAlways P(A|B)=P(B|A)
- DAlways P(A complement) = 1-P(A)
23Probability questions such as “what is the probability of obtaining this even and that event” typically call for the:
- AMultiplication rule
- BDivision rule
- CSubtraction rule
- DAddition rule
Informationen
- Beschreibung
- Most of these are directly from the teacher's sample questions, some are from the book, and some are of my own making.
- Autor
- Brooke Harris
- Sprache
- English
- Aufrufe
- 17
- Schlagwörter
- Bewertung
- Noch keine Bewertungen Bewerte dieses Quiz